2. The East Street Bagel Shop has a weekly demand for 3,500 bagels, which they make at their store fresh every day. They operate 7 days a week. The raw materials and ingredients are supplied by a trusted supplier. On average, the supplier has a 2-day lead-time. They have also decided to keep a safety stock of 250 bagels equivalent to take care of any demand or supply randomness. Keeping this in mind, what is the Reorder Point for the East Street Bagel Shop, in terms of equivalent bagels?
Answer #1

ROP = DEMAND DURING LEAD TIME + SAFETY STOCK

SS = 250

SINCE THE LEAD TIME IS IN DAYS, WE WILL CONVERT THE WEEKLY DEMAND T DAILY

DAILY DEMAND = 3500 / 7 = 500

LEAD TIME = 2 DAYS

REORDER POINT = DEMAND * LT + SS

= 500 * 2 + 250 = 1000 + 250 = 1250